# Fair Overlap Number of Balls (Fair-ONB): A Data-Morphology-based Undersampling Method for Bias Reduction

### v1

@@ Abstract
Given the magnitude of data generation currently, both in quantity and speed, the use of machine learning is increasingly important. When data include protected features that might give rise to discrimination, special care must be taken. Data quality is critical in these cases, as biases in training data can be reflected in classification models. This has devastating consequences and fails to comply with current regulations. Data-Centric Artificial Intelligence proposes dataset modifications to improve its quality. Instance selection via undersampling can foster balanced learning of classes and protected feature values in the classifier. When such undersampling is done close to the decision boundary, the effect on the classifier would be bolstered. This work proposes Fair Overlap Number of Balls (Fair-ONB), an undersampling method that harnesses the data morphology of the different data groups (obtained from the combination of classes and protected feature values) to perform guided undersampling in the areas where they overlap. It employs attributes of the ball coverage of the groups, such as the radius, number of covered instances and density, to select the most suitable areas for undersampling and reduce bias. Results show that the Fair-ONB method reduces bias with low impact on the classifier’s predictive performance.
Keywords Trustworthy Artificial Intelligence  $\cdot$ Data-Centric Artificial Intelligence  $\cdot$ Bias Reduction  $\cdot$ Undersampling  $\cdot$ Data Morphology.
@@ 1 Introduction
The vast amount of data generated in this digital era, ranging from simple sensor measurements to biometric data, makes harnessing the relevant information an impossible task unless some processes are automated. Machine Learning and Artificial Intelligence learn the inner patterns in data, making it possible for computers to predict the behaviour of unseen data.
@@
To learn these patterns and, for instance, the classes in supervised classification problems, the characteristics of each dataset are crucial [10]. Models learned from poor-quality data do not reach high predictive performance and might draw unrealistic conclusions [22]. In particular, the presence of class overlap hinders the model’s estimation of class boundaries, class imbalance can reduce the importance of some classes in the model (which complicates an equable predictive performance) and noisy samples can distort the predictions. Data-Centric Artificial Intelligence is the area of Artificial Intelligence whose aim is to obtain better models by improving data quality [6].
@@
When datasets include personal data or sensitive features that might give rise to discrimination, low data quality can lead to models that include discriminatory bias [12]. This is a critical problem nowadays, given the evolution of regulations in Artificial Intelligence, such as the recently approved European Union Artificial Intelligence Act [^1] Artificial intelligence (AI) act: Council gives final green light to the first worldwide rules on AI https://www.consilium.europa.eu/es/press/press-releases/2024/05/21/artificial-intelligence-ai-act-council-gives-final-green-light-to-the-first-worldwide-rules-on-ai/. Some sensitive features, usually referred to as protected features (for example, race or gender), are prone to causing discrimination amongst groups or classes and, by law, cannot be included in the model’s decision function [^2] AI Act, https://artificialintelligenceact.eu/. However, even if these features are not directly involved in a model, it can still include biases depending on how instances with different values on those protected features are classified (sometimes due to the indirect impact of other highly-correlated features to the protected one that the model does use).
@@
To reduce model bias, a widespread Data-Centric approach is to modify the dataset via sampling [18]. Sampling data of specific groups or classes, either by adding instances of those with fewer data (oversampling) or eliminating some of the over-represented ones (undersampling), can foster an equitable treatment of all groups or classes by the model. When protected features are present, this would mean balancing the model’s performance for instances of various values of the protected features. Specifically, this means balancing the groups formed by the combinations of the class values and protected feature values as, when taken into account separately, not only might bias reduction be harder, but counterproductive situations might arise (for example, if more $gender=female$ samples were added to balance $gender$ but they were added in negative class areas, discrimination against $gender=female$ would increase).
@@
Sampling strategies in the literature follow varied strategies. Most are based on sampling until group balance is reached and, while some strategies do it randomly, others include restrictions or preferences, such as only balancing groups close to the model’s decision boundaries [18]. However, just balancing a priori ratios of protected feature values and classes does not always yield equal performance of the model (particularly, when the distribution of instances with each protected feature value is different).
To properly balance the model’s performance on the different data groups obtained from the combinations of class and protected feature values, using overlap estimation methods could prove beneficial. This way, areas of the data space where sampling would be the most effective towards reducing model bias can be selected. The hypothesis of this work is that, since class or group overlap areas are problematic for model decisions, group morphology can be used to characterise and modify datasets in ways that increase performance and fairness.
@@
Amongst the strategies that locate and measure overlap in datasets, "Overlap Number of Balls" (ONB) was chosen [14]. This technique is based on the coverage of data using balls that only include points from a single group. ONB employs morphology to characterise the different groups in data, as well as to estimate group overlap and classifier performance.
@@
The aim of this work is to leverage the overlap estimation capabilities of ONB and the characterisation of groups (obtained using class and protected feature value combinations) derived from its coverage strategy to improve fairness in classification models according to protected feature values without a noticeable reduction in model performance. To this end, we propose a new preprocessing method via sampling, Fair-ONB, that is based on group morphology and data neighbourhoods. Since undersampling allows for the reduction of both bias and noise and oversampling would only tackle the former, the proposed strategy will perform undersampling.
@@
The Fair-ONB method involves using percentile thresholds on coverage ball attributes, such as ball radius, number of instances covered and density, to select which instances to remove. This new method produces stable results, compared to random sampling strategies, due to using empirical measures to decide which samples to eliminate. Moreover, this instance selection includes the knowledge of the undersampled areas and group size evolution.
@@
The efectiveness of the proposed Fair-ONB method will be tested on multiple of the most broadly-used real datasets in data fairness studies. The a priori behaviour of the different threshold variables for instance filtering is studied, both in terms of predictive performance and fairness (according to disparate impact reduction). The results are compared with FAWOS, a state-of-the-art, neighbourhood-based sampling technique [18].
@@
The rest of this paper is organised as follows. Firstly, Section 2 presents the preliminaries on fairness, discrimination and bias. Then, Section 3 details the proposed Fair-ONB method. Next, Section 4 describes the experimental framework for the evaluation of the Fair-ONB method’s behaviour and its comparison with the state of the art. Immediately after, Section 5 shows the experimental results. Lastly, Section 6 presents the concluding remarks from this manuscript and some future work lines.
@@ 2 Preliminaries on Discrimination and Bias
The existence of discrimination amongst groups of people, that is, the treatment that inherently favours some of them, involves substantial ethical problems and is legally unacceptable. Thus, the use of model bias reduction techniques so that they comply with the current regulations [3] is a currently thriving research area. These biases can be measured in different ways, according to the situation [16, 12]. For example, it might be preferable that an equal rate of acceptance for a grant between men and women but, when assigning a medical procedure, the actual need for such treatment must also be taken into account.
According to these preferences, there are multiple possible approaches in order to determine if a classifier is biased [13]. For example, in some cases the aim is to maintain demographic parity, or statistical parity, which involves the positive prediction being independent on the protected feature; that is, given a prediction Ŷ, demographic parity is measured according to how different the chances of obtaining a prediction Ŷ=1 is for each value of protected feature P. This would create 2 similar metrics:
@@
    $DPE=Pr(\widehat{Y}=1|P=0)-Pr(\widehat{Y}=1|P=1)$ (1)
  • Disparate Impact (DI): it measures the ratio between the probabilities of a positive classification between two protected feature values (Equation 2). In the optimum case (equal chances), its value is 1. Given its extended use to check biases in the USA, this metric is more commonly applied.
    $ID=\frac{Pr(\widehat{Y}=1|P=0)}{Pr(\widehat{Y}=1|P=1)}$ (2)
@@
$IDA=\left\{\begin{array}[]{lr}ID&:ID\leq 1\\ \frac{1}{ID}&:ID>1\end{array}\right.$ (3)
Other extended metrics would involve equality of probability and equality of opportunity, whose foci are the true positive and/or false positive rates being the same for different values of the protected feature. Their associated metrics are measured according to the difference in probabilities for each group, and are named Equal Probability Difference (EPD) and Equal Opportunity Difference (EOD), whose formulae are given in Equations 4 and 5, respectively.
@@
Unless they are based on the same types of probabilities (such as in SPD and DI), optimizing multiple bias metrics simultaneously is generally unfeasible; therefore, each study generally only selects one of them to measure fairness.
Having indicated multiple metrics that measure fairness, knowing how to reduce biases is also important. There are many techniques in that regard and, since this study is focused on the use of sampling, the following Scopus queries were used in order to check the strategies in that area:
TITLE-ABS-KEY ("undersampling")
AND TITLE-ABS-KEY ("fairness")
TITLE-ABS-KEY ("sampling")
AND TITLE-ABS-KEY ("fairness")
AND TITLE-ABS-KEY ("preprocessing")
TITLE-ABS-KEY ("sampling")
AND TITLE-ABS-KEY ("fairness")
AND TITLE-ABS-KEY ("’machine learning’")
After selecting the relevant documents, a total of 30 articles were found (by May of 2024), all of them published between 2018 and 2024.
Their strategies are mainly based on the equal treatment of the subgroups of data defined by the combinations of class and protected feature values (for example, the pair of $Class=1$ and $Gender=F$). This is due to fairness metrics taking both into account simultaneously and, thus, seeking only one class balance or protected feature balance on their own would not be sufficient.
@@ 3 The Fair Overlap Number of Balls method
This Section describes in more detail the functioning of Fair-ONB, a novel undersampling method based on data morphology. This method eliminates samples from groups with favourable bias so that protected groups are treated similarly by the classifier. This is done in a guided manner, harnessing the morphology of the group coverage obtained from the Overlap Number of Balls algorithm and its ball properties in order to select problematic areas of the data space.
Firstly, the ONB algorithm is explained in Section 3.1. Then, some preliminaries on the choices made during the method’s creation are indicated in Section 3.2. Lastly, the structure and inner working of the method are explained in Section 3.3.
@@ 3.1 Overlap Number of Balls
Overlap Number of Balls is a complexity metric, that is, a technique that measures how difficult to classify a dataset is according to its inner characteristics [14]. ONB is based on the Pure Class Cover Catch Digraph (P-CCCD) ball coverage algorithm [11].
Their strategy is simple. Firstly, for each instance in the dataset, the biggest open ball (the interior of the hypersphere of the same radius) that does not cover any points of a different class is generated. Then, for each class, the balls that would cover the most points that are not yet covered by other balls in the coverage set are iteratively chosen as part of said coverage set, until all points of that class are covered. An example of this is shown in Figure 1.
@@
This coverage strategy has very useful implications. Firstly, it estimates the distribution of the classes in the dataset. Secondly, the number of balls that are necessary to cover the dataset give information on both how difficult to classify the dataset is and the degree of overlap amongst classes, which are very useful towards deciding whether preprocessing is necessary in order to improve classifier performance [14]. Thirdly, the ball coverage can estimate class boundaries, either in the base dataset or in the classifier that is learned from it; in the latter case, the morphology of these estimated boundaries can be used to improve the explainability of complex classification models [15].
@@ 3.2 Preliminaries on the Fair-ONB method
Given a dataset whose protected features and class are known, the combination of their values is employed to detect and correct biases. For the Fair-ONB method, each of these combinations forms a group. For example, a dataset with binary protected features “race” and “gender” and a binary class would have the groups shown in Table 1).
@@
At the same time, before employing the method, the existence of a priori biases in the dataset or model is checked using a fairness metric. This allows the selection of which subgroups need to be preprocessed. In this study, said fairness metric is the Disparate Impact (DI, Equation 2), which, as a reminder, is the ratio between the probabilities of obtaining a positive result (Ŷ=1) according to the values in the protected feature.
When no bias is present according to a protected feature, $DI=1$; otherwise, the value is still considered acceptable when it is between 0.8 and 1.25. Figure 2 shows the behaviour of DI as the preferences of the model change (as happens when its training data is modified). It includes the neutral DI value (1) and the positive classification probabilities for the different values of a protected feature.
@@
With this in mind, DI can be checked on the base dataset (with the existing ratios) or on the model (using the predicted classes) for each of the protected features. When $DI>1$, the bias favours the group with protected feature value 0; when $DI<1$, it is biased towards the feature value 1.

### v2

@@ Abstract
One of the key issues regarding classification problems in Trustworthy Artificial Intelligence is ensuring Fairness in the prediction of different classes when protected (sensitive) features are present. Data quality is critical in these cases, as biases in training data can be reflected in machine learning, impacting human lives and failing to comply with current regulations. One strategy to improve data quality and avoid these problems is preprocessing the dataset. Instance selection via undersampling can foster balanced learning of classes and protected feature values. Performing undersampling in class overlap areas close to the decision boundary should bolster the impact on the classifier. This work proposes Fair Overlap Number of Balls (Fair-ONB), an undersampling method that harnesses the data morphology of the different data groups (obtained from the combination of classes and protected feature values) to perform guided undersampling in overlap areas. It employs attributes of the ball coverage of the groups, such as the radius, number of covered instances and density, to select the most suitable areas for undersampling and reduce bias. Results show that the Fair-ONB method improves model Fairness with low impact on the classifier’s predictive performance.
Keywords Trustworthy Artificial Intelligence  $\cdot$ Preprocessing  $\cdot$ Fairness  $\cdot$ Data Morphology.
@@ 1 Introduction
The vast amount of data generated in this digital era [8], ranging from simple sensor measurements to biometric data, makes harnessing the relevant information an impossible task unless some processes are automated. Machine Learning techniques learn the inner patterns in data, making it possible for computers to predict the behaviour of unseen data [1].
@@
To learn these patterns and, for instance, the classes in supervised classification problems, the characteristics of each dataset are crucial [14]. Models learned from poor-quality data do not reach high predictive performance and might draw unrealistic conclusions [26]. In particular, the presence of class overlap hinders the model’s estimation of class boundaries, class imbalance can reduce the importance of some classes in the model (which complicates an equable predictive performance) and noisy samples can distort the predictions.
When datasets include personal data or sensitive features that might give rise to discrimination, low data quality or mistreatment of said features can cause models to include discriminatory bias [16]. This is a critical problem nowadays that has led to auditability requirements and the emergence of Artificial Intelligence regulations, such as the recently approved European Union Regulation 2024/1689111Regulation (EU) 2024/1689 of the European Parliament and of the Council http://data.europa.eu/eli/reg/2024/1689/oj, an evolution of the European Union Artificial Intelligence Act [^2] Artificial Intelligence Act, https://artificialintelligenceact.eu/.
@@
Some sensitive features, usually referred to as protected features (for example, race or gender), are prone to causing discrimination amongst groups or classes and cannot be included in the model’s decision function. However, even if these features are not directly involved in a model, it can still include biases depending on how instances with different values on those protected features are classified (sometimes due to the indirect impact of other highly-correlated features to the protected one that the model does use) [7]. Data and model Fairness are achieved when these possible biases amongst groups are minimised or eliminated.
@@
While Artificial Intelligence often focuses on optimising prediction model parameters, another approach is focusing on data quality, as models learned on high-quality data tend to achieve high performance [14]. To reduce model bias, a widespread approach is to modify the dataset via sampling [22]. Sampling data of specific groups or classes, either by adding instances of those with fewer data (oversampling) or selecting for usage only some from the over-represented ones (undersampling), can foster an equitable treatment of all groups or classes by the model. When protected features are present, this would mean balancing the model’s performance for instances of various values of the protected features. Specifically, this means balancing the groups formed by the combinations of the class values and protected feature values as, when taken into account separately, not only might bias reduction be harder, but counterproductive situations might arise (for example, if more $gender=female$ samples were added to balance $gender$ but they were added in negative class areas, discrimination against $gender=female$ would increase).
@@
Sampling strategies in the literature follow varied strategies. Most are based on sampling until group balance is reached and, while some strategies do it randomly [20], others include restrictions or preferences, such as only balancing groups close to the model’s decision boundaries [22] or using particular clusters [9]. However, just balancing a priori ratios of protected feature values and classes does not always yield equal performance of the model (particularly, when the distribution of instances with each protected feature value is different).
The hypothesis of this work is that, to balance model performance on different data groups, the focus must be on the group overlap areas of the problem, as those would be the sections of the data space were sampling would be the most effective at modifying model decisions and, thus, at improving Fairness. To this aim, using data morphology, which has proven to be a remarkable strategy for the characterisation of groups or classes and the detection of overlap amongst them, would be a well-founded approach.
@@
Amongst the strategies that locate and measure overlap in tabular datasets, the use of "Overlap Number of Balls" (ONB) is proposed [18]. This technique is based on the coverage of data using balls that only include points from a single group. ONB employs morphology to characterise the different groups in data, as well as to estimate group overlap and classifier performance. Furthermore, the information on the boundaries amongst classes or groups provided by the ball coverage can also be harnessed for model explainability [19].
@@
The aim of this work is to leverage the overlap estimation capabilities of ONB and the characterisation of groups (obtained using class and protected feature value combinations) derived from its coverage strategy to improve Fairness in classification models according to protected feature values without a noticeable reduction in model performance. To this end, we propose a preprocessing method via sampling, called Fair Overlap Number of Balls (Fair-ONB), that is based on group morphology and data neighbourhoods. Since undersampling allows for the reduction of both bias and noise (particularly due to using morphology to locate it) and oversampling would only tackle the former, the proposed strategy will perform undersampling.
@@
The Fair-ONB method involves two stages: selecting the undersampling groups, and then using percentile thresholds on coverage ball attributes, such as ball radius, number of instances covered and density, to decide how to undersample. It produces stable results, compared to random sampling strategies [20, 27], due to using empirical measures to decide which samples to eliminate, and the location of balls under the chosen thresholds (and the change in group size that this would entail) can be easily quantified.
@@
The efectiveness of the proposed Fair-ONB method will be tested on several of the most broadly-used real datasets in data Fairness studies. The a priori behaviour of the different threshold variables for instance filtering is studied, both in terms of predictive performance and Fairness. The results are compared with FAWOS, a state-of-the-art, neighbourhood-based sampling technique [22].
@@
The rest of this paper is organised as follows. Firstly, Section 2 presents the preliminaries on Fairness, discrimination and bias. Then, Section 3 details the proposed Fair-ONB method. Next, Section 4 describes the experimental framework for the evaluation of the Fair-ONB method’s behaviour. Section 5 shows the experimental results. Lastly, Section 6 presents the concluding remarks from this manuscript and some future work lines.
@@ 2 Preliminaries on Discrimination and Bias
This Section presents some preliminaries regarding bias in Machine Learning. Section 2.1 introduces important concepts and some of the different notions of Fairness, along with their associated metrics. Section 2.2 shows some of the techniques present in the state of the art for the minimisation of bias, with an emphasis on sampling techniques.
2.1 On Fairness metrics
The existence of discrimination amongst groups of people, that is, the treatment that inherently favours some of them, involves substantial ethical problems and is legally unacceptable [^3] Regulation (EU) 2024/1689 of the European Parliament and of the Council http://data.europa.eu/eli/reg/2024/1689/oj. Fairness in Machine Learning means the absence of bias or discrimination amongst groups of data. This Fairness is checked with respect to protected features, such as gender or race, which are variables in datasets whose unequal treatment would expose the existence of discrimination.
Fairness and bias can be measured in different ways according to the situation, as studied in multiple surveys [16, 17, 3].For example, it might be preferable that an equal rate of acceptance for a grant between men and women but, when assigning a medical procedure, the actual need for such treatment (which might be different according to gender) must also be taken into account.
According to these preferences, there are multiple possible approaches in order to determine if a classifier is biased [17]. For example, in some cases the aim is to maintain demographic parity, also named statistical parity, which involves the positive prediction being independent on the protected feature; that is, given a prediction Ŷ, demographic parity is measured according to how different the chances of obtaining a prediction Ŷ=1 is for each value of protected feature P. This would create 2 similar metrics:
@@ 2.1 On Fairness metrics
    $SPD=Pr(\widehat{Y}=1|P=0)-Pr(\widehat{Y}=1|P=1)$ (1)
  • Disparate Impact (DI): it measures the ratio between the probabilities of a positive classification between two protected feature values (Equation 2). In the optimum case (equal chances), its value is 1. Given its extended use to check biases in the USA, this metric is commonly applied [3].
    $DI=\frac{Pr(\widehat{Y}=1|P=0)}{Pr(\widehat{Y}=1|P=1)}$ (2)
@@
$ADI=\left\{\begin{array}[]{lr}ID&:ID\leq 1\\ \frac{1}{ID}&:ID>1\end{array}\right.$ (3)
Other extended metrics would involve equality of probability (also named equalised odds) and equality of opportunity, whose foci are the true positive and/or false positive rates being the same for different values of the protected feature. Their associated metrics are measured according to the difference in probabilities for each group, and are named Equal Probability Difference (EPD) and Equal Opportunity Difference (EOD), whose formulae are given in Equations 4 and 5, respectively.
@@
Unless they are based on the same types of probabilities (such as in SPD and DI), optimizing multiple bias metrics simultaneously is generally unfeasible [11]; therefore, studies often select one of them to measure Fairness.
2.2 Bias reduction strategies
Having indicated multiple metrics that measure Fairness, knowing how to reduce biases is also important. The use of model bias reduction techniques so that they comply with the current Fairness regulations is a currently thriving research area [7]. Consequently, there are many techniques for bias reduction, from adding Fairness restrictions to the model’s objective function [29], to adversarial training [21], adding weights to data with specific protected feature values [30] or modifying the dataset so it includes less bias before learning the model [22]. Since this study is focused on dataset modifications via sampling, the strategies in that area will now be outlined.
Most sampling strategies are mainly based on the equal treatment of the subgroups of data defined by the combinations of class and protected feature values (for example, the pair of $Class=1$ and $Gender=F$). This is due to Fairness metrics taking both into account simultaneously and, thus, seeking only class balance or protected feature balance on their own would not be sufficient.
@@ 3 The Fair Overlap Number of Balls method
This Section describes in detail the functioning of the Fair-ONB method, the novel undersampling proposal based on data morphology. This method undersamples groups with favourable bias so that protected groups are treated similarly by the classifier. This is done in a guided manner, harnessing the morphology of the group coverage obtained from the Overlap Number of Balls (ONB) algorithm and its ball properties in order to select problematic areas of the data space.
Firstly, the ONB algorithm is explained in Section 3.1. Then, some preliminaries on the choices made during the method’s creation are indicated in Section 3.2. Lastly, the structure and inner working of the method are introduced in Section 3.3.
@@ 3.1 Overlap Number of Balls
ONB is a class boundary estimation method that was initially conceived as a complexity metric, that is, a technique that measures how difficult to classify a dataset is according to its inner characteristics [18]. ONB is based on the Pure Class Cover Catch Digraph (P-CCCD) coverage algorithm [15].
Its strategy is simple. Firstly, for each instance in the dataset, the biggest open ball (the interior of the hypersphere of the same radius) that does not cover any points of a different class is generated. Then, for each class, the balls that would cover the most points that are not yet covered by other balls in the coverage set are iteratively chosen as part of said coverage set, until all points of that class are covered. An example of this is shown in Figure 1.
@@
This coverage strategy has very useful implications.
  • The coverage locates and maps classes in the dataset. This estimation of the class distribution gives information on where it would be normal to find points of each class, and where their presence could be considered unusual or noisy.
  • The number of balls that are necessary to cover the dataset gives information on both how difficult to classify the dataset is and the degree of overlap amongst classes. This information is very useful towards deciding whether preprocessing is necessary in order to improve classifier performance [18].
  • The ball coverage can estimate class boundaries, either in the base dataset or in the classifier that is learnt from it. In the latter case, the morphology of these estimated boundaries can be used to improve the explainability of complex classification models. In particular, Overlap Number of Balls Model-Agnostic Counterfactuals (ONB-MACF) is a counterfactual generator that employs the boundary estimations of ONB to provide counterfactual explanations [19].
@@ 3.2 Preliminaries on the Fair-ONB method
Given a dataset whose protected features and class are known, the combination of their values is employed to detect and correct biases, given their paired involvement in Fairness estimation formulae. For the Fair-ONB method, each of these combinations forms a group. For example, a dataset with binary protected features “race” and “gender” and a binary class would have the groups shown in Table 1).
@@
At the same time, before employing the method, the existence of a priori biases in the dataset or model is checked using a Fairness metric. In this study, said Fairness metric is the Disparate Impact (DI, Equation 2), which, as a reminder, is the ratio between the probabilities of obtaining a positive result (Ŷ=1) according to the values in the protected feature. This first check allows the selection of which subgroups need to be preprocessed.
As shown in Equation 2, when no bias is present according to a protected feature, $DI=1$; otherwise, the value is still considered good when it is between 0.8 and 1.25 [3]. Figure 2 shows the behaviour of DI as the preferences of the model change (as happens when its training data is modified). It includes the neutral DI value (1) and the positive classification probabilities for the different values of a protected feature.
@@
With this in mind, DI can be checked on the base dataset (with the existing ratios, to detect pre-existing bias in the dataset) or on the model (using the predictions given by the classifier, which elucidates whether the model is biased) for each of the protected features. When $DI>1$, the bias favours the group with protected feature value 0; when $DI<1$, it is biased towards the feature value 1.